a number between 10 and 100 is five times the sum of its digits. if 9 be added to it the digits are reversed, find the number.....??????
Answer by chriswen(106)

Let x be the first digit.
Let y be the second digit.
...
5(x+y)=10x+y ...(i)
5x+5y=10x+y
5y-y=10x-5x
4y=5x ...(ii)
...
The 10x+y is because the x digit is in the ten's column and it is worth more.
...
10x+y+9=10y+x ...(iii)
10x-x+9=10y-y
9x+9=9y
9(x+1)=9y
x+1=y ...(iv)
substitute x+1=y into (ii)
4(x+1)=5x ...(v)
4x+4=5x
4=5x-4x
x=4
substitute x=4 into (ii)
4y=5(4)
4y=20
y=20/4
y=5
...
The first digit is 4. The second digit is 5.
The number is 45.